Let a11 x1 + a12 x2 + a13x3 = b1
a21 x1 + a22 x2 + a23 x3 = b2
a31 x1 + a32 x2 + a33 x3 = b3.
Show that if det (A) 6= 0 where det(A) is the determinant of the coefficient matrix,
then x2 =
det(A2)
det(A) where det(A2) is the determinant obtained by replacing the second column of det(A)
by (b1, b2, b3)
T
1
Expert's answer
2020-04-27T18:13:12-0400

Show that if det (A) 6= 0 where det(A) is the determinant of the coefficient matrix

coefficient matrix of the system of equations is

a11 x1 + a12 x2 + a13x3 = b1

a21 x1 + a22 x2 + a23 x3 = b2

a31 x1 + a32 x2 + a33 x3 = b3


then the coefficient matrix A is

A= "\begin{bmatrix}
 a11& a12&a13\\
 a21& a22&a23\\
a31&a32&a33
\end{bmatrix}"

det(A)6=[a11(a22a33-a32a23)-a12(a21a33-a32a23)+a13(a21a32-a31a22)]6

=[[a11(a22a33-a32a23)-a12(a21a33-a32a23)+a13(a21a32-a31a22]


let Y=(a22a33-a32a23) ; Z=a21a33-a32a23 ; L=a21a32-a31a22


det(A)6=[a11Y-a12Z+a13L]6

det(A)^6=[a11Y-a12Z+a13L]6 which is not equal to 0


for: x2=det(A2)

Since we need to replace the second column by (b1,b2,b3), so we first multiply the second column by x2 (when we multiply any determinant by a scalar, it either gets multiplied with a single row or with a single column). Then used the operation to obtain (b1,b2,b3) in second row which you u can easily observe in my solution

then

from the system of linear equations below

a11 x1 + a12 x2 + a13x3 = b1

a21 x1 + a22 x2 + a23 x3 = b2

a31 x1 + a32 x2 + a33 x3 = b3

then:

"x2det(A)=\begin{vmatrix}
 a11 & x2a12&a13 \\
 a21 & x2a22&a23\\
a31&x2a32&a33
\end{vmatrix}"

taking the operation C21=C2+x1C1+x3C3

we have

"x2det(A)=\begin{vmatrix}
 a11 & a11 x1 + a12 x2 + a13x3&a13 \\
 a21 & a21 x1 + a22 x2 + a23 x3&a23\\
a31&a31 x1 + a32 x2 + a33 x3&a33
\end{vmatrix}"


"x2det(A)=\begin{vmatrix}
 a11 & b1&a13 \\
 a21 & b2&a23\\
a31&b3&a33
\end{vmatrix}=det(A2)"


hence

"x2det(A)=det(A2)\\
then\\
x2=det(A2)/det(A)\\" where det(A) is not equal to zero

if det(A)=1 then x2=det(A2)